# Efficient Image-to-Image Diffusion Classifier for Adversarial Robustness

### v1

@@ Abstract
Diffusion models (DMs) have demonstrated great potential in the field of adversarial robustness, where DM-based defense methods can achieve superior defense capability without adversarial training. However, they all require huge computational costs due to the usage of large-scale pre-trained DMs, making it difficult to conduct full evaluation under strong attacks and compare with traditional CNN-based methods. Simply reducing the network size and timesteps in DMs could significantly harm the image generation quality, which invalidates previous frameworks. To alleviate this issue, we redesign the diffusion framework from generating high-quality images to predicting distinguishable image labels. Specifically, we employ an image translation framework to learn many-to-one mapping from input samples to designed orthogonal image labels. Based on this framework, we introduce an efficient Image-to-Image diffusion classifier with a pruned U-Net structure and reduced diffusion timesteps. Besides the framework, we redesign the optimization objective of DMs to fit the target of image classification, where a new classification loss is incorporated in the DM-based image translation framework to distinguish the generated label from those of other classes. We conduct sufficient evaluations of the proposed classifier under various attacks on popular benchmarks. Extensive experiments show that our method achieves better adversarial robustness with fewer computational costs than DM-based and CNN-based methods. The code is available at https://github.com/hfmei/IDC.
@@ Introduction
Diffusion models (DMs) have achieved excellent performance in high-quality generative tasks (Ho, Jain, and Abbeel 2020; Song et al. 2020; Song, Meng, and Ermon 2020; Dhariwal and Nichol 2021; Rombach et al. 2022), leading to a range of relative applications, such as in-painting (Lugmayr et al. 2022; Zhang et al. 2023), audio or video synthesis (Kong et al. 2021; Ho et al. 2022; Singer et al. 2022), image super-resolution (Saharia et al. 2022b; Li et al. 2022), graph generation (Niu et al. 2020) and image-to-image translation (Saharia et al. 2022a; Sasaki, Willcocks, and Breckon 2021; Li et al. 2023b). In addition to the remarkable generating related capabilities, diffusion models also contribute significantly to enhancing the adversarial robustness of classification (Nie et al. 2022; Zhang, Luo, and Zhang 2023; Chen et al. 2023a; Wang et al. 2022; Lin et al. 2024b).
@@
In this paper, we aim to redesign the diffusion framework for adversarial robustness to eliminate this gap, relieving the pressure on DMs to generate high-quality images, which enables an efficient incorporation of DMs in adversarial robustness. Specifically, we regard classification tasks as image translation tasks achieved by DMs, where the image is translated to the pre-defined labels that are presented by different images. Since the objective is to align images with their corresponding image labels instead of high-quality image generation, the required complexity of DMs in adversarial robustness can be significantly reduced. With this simple framework, we introduce an Image-to-Image Diffusion Classifier (IDC) to effectively and efficiently apply DMs to adversarially robust image classification. In order to guarantee the diversity of pre-defined image labels for classification, we propose to construct orthogonal image labels in pixel space. Then IDC employs an image-to-image translation framework to learn a many-to-one mapping from input images to these orthogonal image labels, and the classification can be achieved by measuring the distances between translated results and all pre-defined image labels. Due to the switch of objectives, we are able to alleviate the computation of DMs in different aspects, including pruning the U-Net structure and reducing the diffusion timesteps. Furthermore, the vanilla optimization of translation frameworks only minimizes the distance between the translated result and its corresponding image labels without considering other classes. Thus, we incorporate a classification loss in the optimization of DM-based image translation to better fit the classification objective. Our introduced IDC achieves superior performance in the trade-offs between model efficiency and adversarial robustness. As shown in Figure 1, we achieve more competitive robustness with fewer model parameters than CNN-based adversarial training methods and much lower FLOPs cost than the DM-based method. We summarize the main contributions as follows:
@@
Applying a pre-trained diffusion model to classification tasks requires additional components or computations. For example, diffusion classifiers (Chen et al. 2023a; Li et al. 2023a) perform inference for each class separately, which leads to a computational overhead that scales proportionally with the number of classes $K$, as shown in Figure 2(a). DiffPure (Nie et al. 2022) requires an additional pre-trained classifier and $t_{L}$ steps of the diffusion process, as shown in Figure 2(b). All the iterative steps in these methods heavily rely on the large-scale U-Net (Ronneberger, Fischer, and Brox 2015), which makes their inference phases much slower than traditional CNN-based classifiers. Given these limitations, we are motivated to alleviate huge computational costs in DMs for adversarial robustness, such as the size of U-Net and the number of diffusion steps. However, reducing computation in diffusion models could significantly influence the image generation quality, which invalidates previous frameworks, such as purification in (Nie et al. 2022) and noise predictor in (Chen et al. 2023a). Thus, we propose to redesign the diffusion framework for adversarial robustness, which relieves the demand for high-quality image generation.
@@
Besides the intra-class in Eq. (10), the inter-class loss of the diffusion classifier remains unexplored. Thus, we propose to design a loss $\mathcal{L}_{inter}$ to push the translated image label $\hat{\bm{y}}_{0}$ away from other classes, as shown in Figure 3(c). Formally, we consider the most confusing class $j$ of image $\bm{x}$ where the image label $\bm{y}^{j}_{0}$ is the most closest one to the translated $\hat{\bm{y}}_{0}$, as $j=\operatorname*{argmin}_{i}\|\hat{\bm{y}}_{0}-\bm{y}^{i}_{0}\|_{1}$ and $j\neq i$. In the framework of translation, the distance between $\hat{\bm{y}}^{i}_{0}$ and the most confusing image label $\bm{y}^{j}_{0}$ is expected to be maximized, which can be formulated in a simplified reverse format of ELBO as:
@@
We conduct experiments on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky, Hinton et al. 2009). We compare our classifier with both CNN-based and DM-based methods, using two metrics: standard accuracy and robust accuracy. Different from the setting of most DM-based methods, which randomly select a test subset for robust evaluation, we evaluate the performance of our IDC on the entire test dataset.
@@
We evaluate our methods with CNN-based methods under five attacks: PGD (Madry et al. 2018), FGSM (Szegedy et al. 2014), MIFGSM (Dong et al. 2017), CW (Carlini and Wagner 2017) and AutoAttack (Croce and Hein 2020). For each attack, $\epsilon$ is set to 8/255, the attack steps of PGD and MIFGSM are set to 20 and 5, and the steps and learning rate of CW are 1000 and 0.01. Following (Lee and Kim 2023; Nie et al. 2022), we evaluate DM-based method under BPDA+EOT attack (Hill, Mitchell, and Zhu 2021) with $\ell_{\infty}$ perturbations and PGD+EOT attack (Athalye, Carlini, and Wagner 2018) while $\epsilon=8/255$, steps are 200 and EOT is set to 20. We set the misclassified clean samples as adversarial samples that have successfully attacked our model.
@@
CIFAR-10.
The upper part of Table 1 shows the performance of the CNN-based defense models and our IDC on CIFAR-10 dataset against $\text{PGD}^{20}$, FGSM, MIFGSM, CW and AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$) where $\text{PGD}^{20}$ denotes the attack steps of PGD is 20. Without AT, both ResNet-18 and WideResNet32 perform near zero accuracy under $\text{PGD}^{20}$, MIFGSM, and AutoAttack. In contrast, IDC maintains an accuracy exceeding 80% on $\text{PGD}^{20}$, MIFGSM, and 59.9% on AutoAttack. Even compared with AT methods, IDC improves an average robust accuracy by 20.9% over ResNet-18 and 17.18% over WideResNet32. Among them, our IDC improves the robust accuracy by 32.6% and 29.7% against $\text{PGD}^{20}$ compared with ResNet-18 and WideResNet32.
CIFAR-100.
We compare the robust performance on CIFAR-100 against $\text{PGD}^{20}$, FGSM, MIFGSM, CW and AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$) on the lower part of Table 1. Our IDC maintains competitive standard accuracy with CNN-based classifiers with AT. As for the robust accuracy, the IDC can improve by 4.7%, 10.7%, 8.5% over ResNet-18 and 1.7%, 6.6%, 4.7% over WideResNet32 against $\text{PGD}^{20}$, FGSM, MIFGSM attack. It also shows competitive performance against CW attack. As for AutoAttack, IDC maintains 17.7% robust accuracy compared with zero effectiveness of CNN-based methods without AT.
@@
We compare our method against BPDA+EOT and PGD+EOT attacks in Table 2. Different from other DM-based defense methods, we conduct a full evaluation of CIFAR-10. Compared with ResNet-18 and WideResNet32, IDC improves the robust accuracy by 31.89%, 28.73% on BPDA+EOT attack and 33.53%, 30.45% on PGD+EOT attack with lower parameters. As for DM-based methods, IDC achieves more competitive performance with significantly reduced computations. Concretely, compared with DiffPure (Nie et al. 2022), we improve the accuracy by 1.57% against BPDA+EOT attacks and 37.86% against PGD+EOT attacks, with only 3.84% of the parameters. Similarly, with only 8.42% of the parameters for RDC (Chen et al. 2023a), we achieve a 9.73% accuracy improvement against BPDA+EOT attacks.
@@
(a) Robustness under different PGD attack steps
(b) Robustness under different PGD attack eps
(c) Ablation of diffusion timesteps
Ablation Study
Effectiveness of proposed modules.
@@
For a better illustration of the reduction of diffusion timesteps, we compile statistics on the standard accuracy and robustness accuracy under PGD and AutoAttack of the IDC in Figure 4(c), with timesteps $L_{s}$ of 4, 8, and 12, respectively. The performance at timesteps $L_{s}=8$ is comparable to that at $L_{s}=4$, but it incurs a substantial increase in computational cost during inference. When $L_{s}=12$, the standard performance decreases by 4.5%, and the adversarial robustness declines by 7.1% and 9.6%, respectively. A possible reason is that our framework reduces the generative complexity of DMs. When there are excessive timesteps, the noise addition process in each timestep is not sufficiently pronounced, inadvertently increasing the learning difficulty for the network.
$\alpha$ Standard Acc Robust Acc
$\text{PGD}^{20}$ AutoAttack
$0.15$ 86.6 78.4 56.0
$[0.01,0.2]$ 87.0 84.8 57.6
$[0.01,0.4]$ 86.5 82.0 61.9
$0.2$ 86.0 84.8 59.9
Impact of loss weight $\bm{\alpha}$.

### v2

@@ Abstract
Diffusion models (DMs) have demonstrated great potential in the field of adversarial robustness, where DM-based defense methods can achieve superior defense capability without adversarial training. However, they all require huge computational costs due to the usage of large-scale pre-trained DMs, making it difficult to conduct full evaluation under strong attacks and compare with traditional CNN-based methods. Simply reducing the network size and timesteps in DMs could significantly harm the image generation quality, which invalidates previous frameworks. To alleviate this issue, we redesign the diffusion framework from generating high-quality images to predicting distinguishable image labels. Specifically, we employ an image translation framework to learn many-to-one mapping from input samples to designed orthogonal image labels. Based on this framework, we introduce an efficient Image-to-Image diffusion classifier with a pruned U-Net structure and reduced diffusion timesteps. Besides the framework, we redesign the optimization objective of DMs to fit the target of image classification, where a new classification loss is incorporated in the DM-based image translation framework to distinguish the generated label from those of other classes. We conduct sufficient evaluations of the proposed classifier under various attacks on popular benchmarks. Extensive experiments show that our method achieves better adversarial robustness with fewer computational costs than DM-based and CNN-based methods. The code is available at https://github.com/hfmei/IDC
@@ Introduction
Diffusion models (DMs) have achieved excellent performance in high-quality generative tasks (Ho, Jain, and Abbeel 2020; Song et al. 2020; Song, Meng, and Ermon 2020; Dhariwal and Nichol 2021; Rombach et al. 2022; Ma et al. 2024a, b, c), leading to a range of relative applications, such as in-painting (Lugmayr et al. 2022; Zhang et al. 2023), audio or video synthesis (Ho et al. 2022; Singer et al. 2022; Wang et al. 2024; Chen et al. 2024), image super-resolution (Saharia et al. 2022b; Li et al. 2022) and image-to-image translation (Saharia et al. 2022a; Sasaki, Willcocks, and Breckon 2021; Li et al. 2023b).
@@
In this paper, we aim to redesign the diffusion framework for adversarial robustness to eliminate this gap, relieving the pressure on DMs to generate high-quality images, which enables an efficient incorporation of DMs in adversarial robustness. Specifically, we regard classification tasks as image translation tasks achieved by DMs, where the image is translated to the pre-defined labels that are presented by different images. Since the objective is to align images with their corresponding image labels instead of high-quality image generation, the required complexity of DMs in adversarial robustness can be significantly reduced. With this simple framework, we introduce an Image-to-Image Diffusion Classifier (IDC) to effectively and efficiently apply DMs to adversarially robust image classification. In order to guarantee the diversity of pre-defined image labels for classification, we propose to construct orthogonal image labels in pixel space. Then IDC employs an image-to-image translation framework to learn a many-to-one mapping from input images to these orthogonal image labels, and the classification can be achieved by measuring the distances between translated results and all pre-defined image labels. Due to the switch of objectives, we are able to alleviate the computation of DMs in different aspects, including pruning the U-Net structure and reducing the diffusion timesteps. Furthermore, the vanilla optimization of translation frameworks only minimizes the distance between the translated result and its corresponding image labels without considering other classes. Thus, we incorporate a classification loss in the optimization of DM-based image translation to better fit the classification objective. Our introduced IDC achieves superior performance in the trade-offs between model efficiency and adversarial robustness. As shown in Figure 1, we achieve more competitive robustness with fewer model parameters than CNN-based adversarial training methods and much lower FLOPs cost than the DM-based method.
We summarize the main contributions as follows:
@@
Applying a pre-trained diffusion model to classification tasks requires additional components or computations. For example, diffusion classifiers (Chen et al. 2023a; Li et al. 2023a) perform inference for each class separately, which leads to a computational overhead that scales proportionally with the number of classes $K$, as shown in Figure 2(a). DiffPure (Nie et al. 2022) requires an additional pre-trained classifier and $t_{L}$ steps of the diffusion process, as shown in Figure 2(b). All the iterative steps in these methods heavily rely on the large-scale U-Net (Ronneberger, Fischer, and Brox 2015), which makes their inference phases much slower than traditional CNN-based classifiers.
Given these limitations, we are motivated to alleviate huge computational costs in DMs for adversarial robustness, such as the size of U-Net and the number of diffusion steps. However, reducing computation in diffusion models could significantly influence the image generation quality, which invalidates previous frameworks, such as purification in (Nie et al. 2022) and noise predictor in (Chen et al. 2023a). Thus, we propose to redesign the diffusion framework for adversarial robustness, which relieves the demand for high-quality image generation.
@@
Besides the intra-class in Eq. (10), the inter-class loss of the diffusion classifier remains unexplored. Thus, we propose to design a loss $\mathcal{L}_{inter}$ to push the translated image label $\hat{\bm{y}}^{i}_{0}$ away from other classes, as shown in Figure 3(c). Formally, we consider the most confusing class $j$ of image $\bm{x}$ where the image label $\bm{y}^{j}_{0}$ is the most closest one to the translated $\hat{\bm{y}}^{i}_{0}$, as $j=\operatorname*{argmin}_{j}\|\hat{\bm{y}}^{i}_{0}-\bm{y}^{j}_{0}\|_{1}$ and $j\neq i$. In the framework of translation, the distance between $\hat{\bm{y}}^{i}_{0}$ and the most confusing image label $\bm{y}^{j}_{0}$ is expected to be maximized, which can be formulated in a simplified reverse format of ELBO as:
@@
We conduct experiments on the CIFAR-10, CIFAR-100 datasets (Krizhevsky, Hinton et al. 2009) and Tiny-ImageNet (Deng et al. 2009). We compare our classifier with CNN-based and DM-based methods, using two metrics: standard accuracy and robust accuracy. Different from the setting of most DM-based methods, which randomly select a test subset for robust evaluation, we evaluate the performance of our IDC on the entire test dataset.
@@
We evaluate our methods with CNN-based methods under five attacks: PGD (Madry et al. 2018), FGSM (Szegedy et al. 2014), MIFGSM (Dong et al. 2017), CW (Carlini and Wagner 2017) and AutoAttack (Croce and Hein 2020). For each attack, $\epsilon$ is set to 8/255, the attack steps of PGD and MIFGSM are set to 20 and 5, and the steps and learning rate of CW are 1000 and 0.01. Following (Lee and Kim 2023; Nie et al. 2022), we evaluate DM-based method under BPDA+EOT attack (Hill, Mitchell, and Zhu 2021) with $\ell_{\infty}$ perturbations and PGD+EOT attack (Athalye, Carlini, and Wagner 2018) while $\epsilon=8/255$, steps are 200 and EOT is set to 20.
@@
Method AT Natural $\text{PGD}^{20}$ FGSM CW AutoAttack
ResNet-18 38.06 20.21 21.44 32.52 16.91
44.40 0.26 5.14 12.23 0.00
WideResNet32 44.63 24.21 25.88 38.41 20.93
46.80 0.01 0.84 12.95 0.0
IDC(Ours) 37.65 16.03 22.84 34.33 13.48$\dagger$
Table 2: The adversarial robustness comparison with CNN-based defense methods on Tiny-ImageNet.
@@
CIFAR-10.
The upper part of Table 1 shows the performance of the CNN-based defense models and our IDC on CIFAR-10 dataset against $\text{PGD}^{20}$, FGSM, MIFGSM, CW and AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$) where $\text{PGD}^{20}$ denotes the attack steps of PGD is 20. IDC maintains an accuracy exceeding 80% on $\text{PGD}^{20}$, MIFGSM, and 59.9% on AutoAttack while CNN-based methods without AT perform near zero accuracy. Even compared with AT methods, IDC improves an average robust accuracy by 20.9% over ResNet-18 and 17.18% over WideResNet32. Among them, our IDC improves the robust accuracy by 32.6% and 29.7% against $\text{PGD}^{20}$ compared with ResNet-18 and WideResNet32.
CIFAR-100.
We compare the robust performance on CIFAR-100 against $\text{PGD}^{20}$, FGSM, MIFGSM, CW and AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$) on the lower part of Table 1. Our IDC maintains competitive standard accuracy with CNN-based classifiers with AT. As for the robust accuracy, the IDC can improve by 4.7%, 10.7%, 8.5% over ResNet-18 and 1.7%, 6.6%, 4.7% over WideResNet32 against $\text{PGD}^{20}$, FGSM, MIFGSM attack. It also shows competitive performance against CW attack and AutoAttack.
@@
Tiny-ImageNet.
Compared to AT methods, IDC performs better than ResNet-18 and slightly less than WideResNet32. However, compared to methods without AT, IDC has significant performance advantages. We will consider improving the robustness of IDC in complex datasets in the future.
(a) Robustness under different PGD attack steps
(b) Robustness under different PGD attack eps
(c) Ablation of diffusion timesteps
@@
We compare our method against BPDA+EOT and PGD+EOT attacks in Table 3. Different from other DM-based defense methods, we conduct a full evaluation of CIFAR-10. Compared with ResNet-18 and WideResNet32, IDC improves the robust accuracy by 31.89%, 28.73% on BPDA+EOT attack and 33.53%, 30.45% on PGD+EOT attack with lower parameters. As for DM-based methods, IDC achieves more competitive performance with significantly reduced computations. Concretely, compared with DiffPure (Nie et al. 2022), we improve the accuracy by 1.57% against BPDA+EOT attacks and 37.86% against PGD+EOT attacks, with only 3.84% of the parameters. Compared with novel approaches RDC, MimicDiffusion (Song et al. 2024) and ContrastDiffPure (Bai et al. 2024), IDC also shows competitive performance with exceptionally lower parameters.
@@
For a better illustration of the reduction of diffusion timesteps, we compile statistics on the standard accuracy and robustness accuracy under PGD and AutoAttack of the IDC in Figure 4(c), with timesteps $L_{s}$ of 4, 8, and 12, respectively. The performance at timesteps $L_{s}=8$ is comparable to that at $L_{s}=4$, but it incurs a substantial increase in computational cost during inference. When $L_{s}=12$, the standard performance decreases by 4.5%, and the adversarial robustness declines by 7.1% and 9.6%, respectively. A possible reason is that when there are excessive timesteps, the noise addition process in each timestep is not sufficiently pronounced, inadvertently increasing the learning difficulty for the network.
@@ C.2 Impact of different pruning modes
$\alpha$ Standard Acc Robust Acc
$\text{PGD}^{20}$ AutoAttack
$0.15$ 86.6 78.4 56.0
$[0.01,0.2]$ 87.0 84.8 57.6
$[0.01,0.4]$ 86.5 82.0 61.9
$0.2$ 86.0 84.8 59.9
C.3 Impact of loss weight $\bm{\alpha}$.